# Regret Analysis: a control perspective

### v1

@@ 1 Introduction
Machine learning and adaptive control have many intersections. As pointed out recently in gaudio2019connections, they have common: update laws, robustness modifications (leaky integrator = regularization), both can use projection, have early stopping procedures, adaptive gain/step-size methods, higher order update laws, and the list goes on. The similarities are even more striking in the area of online learning. The purpose of this work is to discuss similarities and differences taken by the adaptive community as compared to the the online learning community both in terms of performance goals and algorithms. We now discuss the origins of optimal control, adaptive control, and adaptive filtering (as they are pertinent to this discussion). [^1] This short literature review is not meant to be exhaustive This history is important, because there was a divergence in original literature that seems to have carried on to the present.
@@
This brief historical review begins with the modernization of control theory in the 1960s by Rudolf Kálmán as he mathematized the state space approach to linear dynamical systems [doi:10.1137/0301010], introduced algebraic tools to the control field [Kalman1503], and offered his solution to optimal filtering [10.1115/1.3662552]. Several scientist involved in the development of optimal control (for known dynamics) were also beginning to think about optimal control of uncertain systems. Kálmán’s own work in control of uncertain systems can be found in [kalman1958design] where he “examines the problem of building a machine which adjusts itself automatically to control an arbitrary dynamic process.” It was originally thought that by the Certainty Equivalence (CE) principle one could iteratively estimate model parameters and then assuming those parameters were ground truth, derive optimal control gains [simon1956dynamic]. Proving that such a scheme resulted in stable dynamics was challenging and CE as a principal didn’t gain much traction until some 60 years later [mania2019certainty].
@@
At the same time others were trying an entirely different approach, which would later become known as model reference adaptive control. In model reference adaptive control a desired reference dynamical system (or reference trajectory) is defined and then control gains are tuned online (in real-time) so as to drive the uncertain dynamical system toward the reference system. The first use of model reference adaptive control is credited to whitaker1959adaptive where control parameters were trained in real time using gradient descent. Gradient descent within the context of adaptive control at that time was called the “MIT rule”. [^2] The MIT rule in its original form was not a stable algorithm for gradient descent learning with dynamical systems. Interestingly, just a year later widrow1960adaptive showed how to adaptively train a pattern classification machine, using gradient descent as well. The use of gradient descent within the latter context was called the Widrow-Hoff Algorithm, which would later be referred to as the Least Mean Square (LMS) Algorithm [widrow1985adaptive, haykin2002adaptive]. These two publications, just one year apart, using the same principle, kick started model reference adaptive control and machine learning.
@@
In model reference adaptive control one has two objectives 1) prove that all time varying parameters/states of the system are bounded, and 2) that the instantaneous error between the adaptively controlled system and a reference system converges to zero over time (or at least a compact set). Importantly, results are derived for arbitrary initial conditions in terms of the state of the plant, and even the initial condition of the plant and controller can be such that the closed loop system is initially unstable. Unfortunately, at the onset of adaptive control heuristics for online gain tuning were used and denoted as “adaptive” and in 1967 an experimental X-15 aircraft crashed with such an “adaptive” controller [dydek2010adaptive]. Following the crash several decades of work in robust adaptive control followed, resulting in the mature field it is now [narendra1989stable, ioannou1995robust], where it is even back to flying [lavretsky2013robust, lavretsky2019design].
@@
The origins of regret analysis for online learning are credited to blackwell1956 and hannan1957approximation where they were interested in optimizing repeated play games. Taking the one player verses nature analogy from blackwell1956, at each time $t$ a player chooses its best strategy $\hat{\theta}_{t}$ based on past information, and then nature returns the loss $\ell_{t}(\hat{\theta}_{t})$ after that strategy is played. Note that nature’s loss function, $\ell_{t}$, is time varying (or could be adversarially chosen by a second player at each step). Therefore it is not unreasonable to have accrued losses over time that grow unbounded. Therefore, when considering games like this, regret analysis was proposed so that learning strategies could be compared to the single best fixed strategy in hindsight. Regret analysis has been a part of the online learning community since with regret analysis also appearing in seminal work by valiant1984theory and Littlestone1988 when trying to learn Boolean functions online. For a rigorous treatment of online learning and online convex optimization see cesa2006prediction, MAL-018, OPT-013. A brief history of seminal work along the way (that informs our understanding of the field) is in the following footnote. [^3] Brief history of regret analysis for online learning of convex functions: quadratic cost with linear functions [cesa1996worst], convex cost with regression, [cesa1999analysis] general convex optimization [Zinkevich:2003:OCP:3041838.3041955], logarithmic regret for strongly convex functions [hazan2007logarithmic], and adaptive learning rates [hazan2008adaptive, duchi2011adaptive]
@@ 2 The basics of regret analysis for online convex optimization
The learning rates and corresponding regret bounds for online convex optimization are as follows [hazan2019lecture], [levy17]:
@@
(following [hazan2019lecture]) We begin by deriving a bound between the one time point step ahead parameter estimate $\hat{\theta}_{t+1}$ and the optimal fixed parameter value in hindsight, $\bar{\theta}$, defined in (1) as follows
@@
We will reiterate again that at the crux of the proof that online convex optimization with gradient descent attains sub-linear regret is being able to balance the two terms $\frac{1}{\eta_{T}}$ and $\sum\eta_{t}$. Another frequently used learning rate that can perform this balancing act is $\eta_{t}\propto\frac{1}{\sqrt{\sum_{\tau=1}^{t}\left\lVert\nabla\ell_{\tau}\right\rVert^{2}}}$. Methods deploying this type of learning rate have been referred to as adaptive subgradient algorithms and have become wildly popular for training deep neural networks ([duchi2011adaptive]). With the following lemma and corollary one is able to show $\sum_{t=1}^{T}\frac{\left\lVert\nabla\ell_{t}\right\rVert^{2}}{\sqrt{\sum_{\tau=1}^{t}\left\lVert\nabla\ell_{\tau}\right\rVert^{2}}}=O(\sqrt{T})$ which in turn can be used to prove $O(\sqrt{T})$ regret.
@@
([DBLP:journals/corr/abs-1002-4908]) Let $\{b_{t}\}$ be a sequence of elements in $\mathbb{R}_{\geq 0}$ indexed by $t\in\mathbb{N}_{>0}$. Then the following bound holds $\sum_{t=1}^{T}\frac{b_{t}}{\sqrt{\sum_{\tau=1}^{t}{b_{\tau}}}}\leq 2\sqrt{\sum_{t=1}^{T}b_{t}}$.
@@
3 Online (real-time streaming) regression using analysis from the control community
@@
(following goodwin1980discrete) Define a Lyapunov candidate $V_{t}=\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$ (note that this is not the loss function at time $t$, $\tfrac{1}{2}e_{t}^{2}$). The one time discrete difference of the candidate function is as follows $\Delta V_{t}\triangleq V_{t+1}-V_{t}=\tilde{\theta}_{t+1}^{\mathsf{T}}\tilde{\theta}_{t+1}-\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$. Substituting the update law from (2) results in
@@
The crux to this proof is ensuring that we have a learning rate that is smaller than the squared 2-norm of the features at any instance in time. This kind of scaling also appears in what is called Normalized LMS, but we were unable to find any of the literature stemming from [widrow1985adaptive, haykin2002adaptive] that had this style of proof with these strong guarantees. It is also important to note that we did not prove that $\hat{\theta}_{t}$ tends to $\theta_{*}$ but we were still able to prove that $\hat{y}_{t}$ asymptotically converges to $y_{t}$. In order to prove that $\hat{\theta}_{t}$ converges to $\theta_{*}$ extra conditions are required on the regressor vector $x_{t}$, namely that $x_{t}$ is Persistently Exciting (PE) ([doi:10.1137/15M1047805]). [^6] The vector $x_{t}\in\mathbb{R}^{n}$ is PE if there exists $\beta,T>0$ such that $\sum_{\tau=t}^{T+t}x_{\tau}x_{\tau}^{\mathsf{T}}\succeq\beta I\quad\forall\ t.$ Having the regressor vector PE is a necessary and sufficient condition for exponential stability in the large for the point $\tilde{\theta}_{t}=0.$
We concede that even with such strong guarantees the model is “simple” by ML standards and the reference model is not adversarial either. Within the adaptive control literature the closest things we have to adversarial types of proofs are when stability is proved in the presence of unmodelled dynamics or bounded disturbances. This line of work has lead to several robustness modifications [ioannou1995robust, narendra1989stable] and we present a flavor of that kind of model and learning rule modification here now.
@@
where $d_{t}$ is now a bounded time varying disturbance and $\theta_{*}\in\Theta$ for some known bounded convex set $\Theta$. In order to prove that we can stably adapt to the disturbed reference we will need to modify the standard gradient descent law and thus we will be using the parameter update in (4) where the $\sigma$-modification ([ioannou1995robust]) term $\sigma_{t}$ is defined as
@@
(following [ioannou1986robust]) Define a Lyapunov candidate $V_{t}=\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$ as before and $\Delta V_{t}\triangleq V_{t+1}-V_{t}=\tilde{\theta}_{t+1}^{\mathsf{T}}\tilde{\theta}_{t+1}-\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$. Evaluating $\Delta V_{t}$ with the gradient update in (4) and after completing the square the following bound holds
@@ 4 Contrasting the two approaches and the new paradigm of Online Adaptive Control
Now this all straight forward but in the previous section we emphasized that all regret based update laws had this interesting property of vanishing learning rates. Here, none of the parameters are updated with gradient descent and thus there is no learning rate. We do still however have a hyper-parameter in this paradigm that directly controls how much learning is occurring and that is the user specified exploration signal variance $\sigma_{\xi,t}^{2}$. In order to be regret optimal the exploration signal must vanish, with one popular choice being[mania2019certainty]
@@
This was just one example. There are other flavors of regret based optimal control whereby the exploration signal does not decay over time, but instead the control parameters themselves are updated via gradient descent with a time-varying and asymptomatically decreasing learning rate [simchowitz2020improper, hazan2020nonstochastic]. Either directly or indirectly to be regret optimal the rate at which you learn must decrease over time.
@@ 5 Closing
We close with a question. Are regret optimal control strategies a good idea? Any time an exploration signal is active it is unlikely that the controller will be in a deployed state. By deployed state we mean when the controller is needed to actively regulate the system (or track a reference command) in a safety critical or high-performance setting, like an aircraft mid-flight. By minimizing regret in its current paradigm OAC is trying to do two things at once (learn and regulate) which are individually being performed in a suboptimal fashion. If you accept the above claim regarding exploration and a “non-deployed state” then any learning based control system that requires an exploration signal might better be described as an episodic strategy with a learning phase (or state) and a control phase [^7] The term “episodic strategy” was suggested to the authors as the appropriate terminology for this style of learning and control but we cant remember who suggested this to us. Please reach out and we will credit you in the acknowledgments. In an episodic regime you would want to explore as much as possible during the learning phase and not have a decaying exploration signal. We also want to note that while “online” within the ML community is synonymous with regret analysis, this is not the case within the broader engineering community. If one were hearing the words “online adaptive control” for the first time one might assume that you are referring to a real-time control strategy that adapts while controlling. If you accept our above claim about exploration signals and episodic learning then OAC would not be ”adaptive” while in the control phase.

### v2

@@ 1 Introduction
Machine learning and adaptive control have many intersections. As pointed out recently in Gaudio et al. [7], they have common: update laws, robustness modifications (leaky integrator = regularization), both can use projection, have early stopping procedures, adaptive gain/step-size methods, higher order update laws, and the list goes on. The similarities are even more striking in the area of online learning. The purpose of this work is to discuss similarities and differences taken by the adaptive community as compared to the the online learning community both in terms of performance goals and algorithms. We now discuss the origins of optimal control, adaptive control, and adaptive filtering (as they are pertinent to this discussion). [^1] This short literature review is not meant to be exhaustive This history is important, because there was a divergence in original literature that seems to have carried on to the present.
@@
This brief historical review begins with the modernization of control theory in the 1960s by Rudolf Kálmán as he mathematized the state space approach to linear dynamical systems [21], introduced algebraic tools to the control field [22], and offered his solution to optimal filtering [20]. Several scientist involved in the development of optimal control (for known dynamics) were also beginning to think about optimal control of uncertain systems. Kálmán’s own work in control of uncertain systems can be found in [19] where he “examines the problem of building a machine which adjusts itself automatically to control an arbitrary dynamic process.” It was originally thought that by the Certainty Equivalence (CE) principle one could iteratively estimate model parameters and then assuming those parameters were ground truth, derive optimal control gains [32]. Proving that such a scheme resulted in stable dynamics was challenging and CE as a principal didn’t gain much traction until some 60 years later [27].
@@
At the same time others were trying an entirely different approach, which would later become known as model reference adaptive control. In model reference adaptive control a desired reference dynamical system (or reference trajectory) is defined and then control gains are tuned online (in real-time) so as to drive the uncertain dynamical system toward the reference system. The first use of model reference adaptive control is credited to Whitaker [34] where control parameters were trained in real time using gradient descent. Gradient descent within the context of adaptive control at that time was called the “MIT rule”. [^2] The MIT rule in its original form was not a stable algorithm for gradient descent learning with dynamical systems. Interestingly, just a year later Widrow and Hoff [36] showed how to adaptively train a pattern classification machine, using gradient descent as well. The use of gradient descent within the latter context was called the Widrow-Hoff Algorithm, which would later be referred to as the Least Mean Square (LMS) Algorithm [35, 10]. These two publications, just one year apart, using the same principle, kick started model reference adaptive control and machine learning.
@@
In model reference adaptive control one has two objectives 1) prove that all time varying parameters/states of the system are bounded, and 2) that the instantaneous error between the adaptively controlled system and a reference system converges to zero over time (or at least a compact set). Importantly, results are derived for arbitrary initial conditions in terms of the state of the plant, and even the initial condition of the plant and controller can be such that the closed loop system is initially unstable. Unfortunately, at the onset of adaptive control heuristics for online gain tuning were used and denoted as “adaptive” and in 1967 an experimental X-15 aircraft crashed with such an “adaptive” controller [6]. Following the crash several decades of work in robust adaptive control followed, resulting in the mature field it is now [29, 17], where it is even back to flying [24, 23].
@@
The origins of regret analysis for online learning are credited to Blackwell [1] and Hannan [9] where they were interested in optimizing repeated play games. Taking the one player verses nature analogy from Blackwell [1], at each time $t$ a player chooses its best strategy $\hat{\theta}_{t}$ based on past information, and then nature returns the loss $\ell_{t}(\hat{\theta}_{t})$ after that strategy is played. Note that nature’s loss function, $\ell_{t}$, is time varying (or could be adversarially chosen by a second player at each step). Therefore it is not unreasonable to have accrued losses over time that grow unbounded. Therefore, when considering games like this, regret analysis was proposed so that learning strategies could be compared to the single best fixed strategy in hindsight. Regret analysis has been a part of the online learning community since with regret analysis also appearing in seminal work by Valiant [33] and Littlestone [26] when trying to learn Boolean functions online. For a rigorous treatment of online learning and online convex optimization see Cesa-Bianchi and Lugosi [3], Shalev-Shwartz [30], Hazan [11]. A brief history of seminal work along the way (that informs our understanding of the field) is in the following footnote. [^3] Brief history of regret analysis for online learning of convex functions: quadratic cost with linear functions [4], convex cost with regression, [2] general convex optimization [37], logarithmic regret for strongly convex functions [13], and adaptive learning rates [14, 5]
@@ 2 The basics of regret analysis for online convex optimization
The learning rates and corresponding regret bounds for online convex optimization are as follows [12], [25]:
@@
(following [12]) We begin by deriving a bound between the one time point step ahead parameter estimate $\hat{\theta}_{t+1}$ and the optimal fixed parameter value in hindsight, $\bar{\theta}$, defined in (1) as follows
@@
We will reiterate again that at the crux of the proof that online convex optimization with gradient descent attains sub-linear regret is being able to balance the two terms $\frac{1}{\eta_{T}}$ and $\sum\eta_{t}$. Another frequently used learning rate that can perform this balancing act is $\eta_{t}\propto\frac{1}{\sqrt{\sum_{\tau=1}^{t}\left\lVert\nabla\ell_{\tau}\right\rVert^{2}}}$. Methods deploying this type of learning rate have been referred to as adaptive subgradient algorithms and have become wildly popular for training deep neural networks ([5]). With the following lemma and corollary one is able to show $\sum_{t=1}^{T}\frac{\left\lVert\nabla\ell_{t}\right\rVert^{2}}{\sqrt{\sum_{\tau=1}^{t}\left\lVert\nabla\ell_{\tau}\right\rVert^{2}}}=O(\sqrt{T})$ which in turn can be used to prove $O(\sqrt{T})$ regret.
@@
([28]) Let $\{b_{t}\}$ be a sequence of elements in $\mathbb{R}_{\geq 0}$ indexed by $t\in\mathbb{N}_{>0}$. Then the following bound holds $\sum_{t=1}^{T}\frac{b_{t}}{\sqrt{\sum_{\tau=1}^{t}{b_{\tau}}}}\leq 2\sqrt{\sum_{t=1}^{T}b_{t}}$.
@@
3 Online regression using analysis from the control community
@@
(following Goodwin et al. [8]) Define a Lyapunov candidate $V_{t}=\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$ (note that this is not the loss function at time $t$, $\tfrac{1}{2}e_{t}^{2}$). The one time discrete difference of the candidate function is as follows $\Delta V_{t}\triangleq V_{t+1}-V_{t}=\tilde{\theta}_{t+1}^{\mathsf{T}}\tilde{\theta}_{t+1}-\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$. Substituting the update law from (2) results in
@@
The crux to this proof is ensuring that we have a learning rate that is smaller than the squared 2-norm of the features at any instance in time. This kind of scaling also appears in what is called Normalized LMS, but we were unable to find any of the literature stemming from [35, 10] that had this style of proof with these strong guarantees. It is also important to note that we did not prove that $\hat{\theta}_{t}$ tends to $\theta_{*}$ but we were still able to prove that $\hat{y}_{t}$ asymptotically converges to $y_{t}$. In order to prove that $\hat{\theta}_{t}$ converges to $\theta_{*}$ extra conditions are required on the regressor vector $x_{t}$, namely that $x_{t}$ is Persistently Exciting (PE) ([18]). [^6] The vector $x_{t}\in\mathbb{R}^{n}$ is PE if there exists $\beta,T>0$ such that $\sum_{\tau=t}^{T+t}x_{\tau}x_{\tau}^{\mathsf{T}}\succeq\beta I\quad\forall\ t.$ Having the regressor vector PE is a necessary and sufficient condition for exponential stability in the large for the point $\tilde{\theta}_{t}=0.$
We concede that even with such strong guarantees the model is “simple” by ML standards and the reference model is not adversarial either. Within the adaptive control literature the closest things we have to adversarial types of proofs are when stability is proved in the presence of unmodelled dynamics or bounded disturbances. This line of work has lead to several robustness modifications [17, 29] and we present a flavor of that kind of model and learning rule modification here now.
@@
where $d_{t}$ is now a bounded time varying disturbance and $\theta_{*}\in\Theta$ for some known bounded convex set $\Theta$. In order to prove that we can stably adapt to the disturbed reference we will need to modify the standard gradient descent law and thus we will be using the parameter update in (4) where the $\sigma$-modification ([17]) term $\sigma_{t}$ is defined as
@@
(following [16]) Define a Lyapunov candidate $V_{t}=\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$ as before and $\Delta V_{t}\triangleq V_{t+1}-V_{t}=\tilde{\theta}_{t+1}^{\mathsf{T}}\tilde{\theta}_{t+1}-\tilde{\theta}_{t}^{\mathsf{T}}\tilde{\theta}_{t}$. Evaluating $\Delta V_{t}$ with the gradient update in (4) and after completing the square the following bound holds
@@ 4 Contrasting the two approaches and the new paradigm of Online Adaptive Control
Now this all straight forward but in the previous section we emphasized that all regret based update laws had this interesting property of vanishing learning rates. Here, none of the parameters are updated with gradient descent and thus there is no learning rate. We do still however have a hyper-parameter in this paradigm that directly controls how much learning is occurring and that is the user specified exploration signal variance $\sigma_{\xi,t}^{2}$. In order to be regret optimal the exploration signal must vanish, with one popular choice being[27]
@@
This was just one example. There are other flavors of regret based optimal control whereby the exploration signal does not decay over time, but instead the control parameters themselves are updated via gradient descent with a time-varying and asymptomatically decreasing learning rate [31, 15]. Either directly or indirectly to be regret optimal the rate at which you learn must decrease over time.
@@ 5 Closing
We close with a question. Are regret optimal control strategies deployable? Any time an exploration signal is active it is unlikely that the controller will be in a deployed state. By deployed state we mean when the controller is needed to actively regulate the system (or track a reference command) in a safety critical or high-performance setting, like an aircraft mid-flight. By minimizing regret in its current paradigm OAC is trying to do two things at once (learn and regulate) which are individually being performed in a suboptimal fashion. If you accept the above claim regarding exploration and a “non-deployed state” then any learning based control system that requires an exploration signal would really only be utilized in an episodic paradigm with a learning phase (identification phase/state) and a control phase (deployed phase/state) [^7] The term “episodic strategy” was suggested to the authors as the appropriate terminology for this style of learning and control but we cant remember who suggested this to us. Please reach out and we will credit you in the acknowledgments. With an episodic paradigm you would want to explore as much as possible during the learning phase and not have a decaying exploration signal. We also want to note that while “online” within the ML community is synonymous with regret analysis, this is not the case within the broader engineering community. If one were hearing the words “online adaptive control” for the first time one might assume that you are referring to a real-time control strategy that adapts while controlling. If you accept our above claim about exploration signals and episodic learning then OAC would not be “adapting” in real-time while it was in a deployed state.